# Propensity Score Methods for Local Test Score Equating: Stratification and Inverse Probability Weighting

### v1

@@ 3.2.2 Assumptions and Properties
The IPW-stratified approach offers several advantages over global weighting. First, it allows for heterogeneous equating relationships across different regions of the propensity score distribution, capturing potential effect modification by covariates. Second, by conducting weight trimming within strata, we can better control the influence of extreme weights while preserving local equating relationships. However, this flexibility comes at the cost of requiring sufficient sample sizes within each stratum to ensure stable estimation of the local equating functions.
@@ 4.1 Equipercentile Equating
While our primary focus has been on linear equating, the framework can be extended to equipercentile equating. Equipercentile equating assumes that scores on different forms of a test should have the same percentile ranks in the populations of examinees taking each form. Let $F_{X}$ and $F_{Y}$ denote the cumulative distribution functions of scores on forms $X$ and $Y$, respectively. The equipercentile equating transformation is defined as:
$\varphi_{EP}(y)=F_{X}^{-1}(F_{Y}(y))$
To implement this within our IPW framework for example, we can estimate $F_{X}$ and $F_{Y}$ using weighted empirical distribution functions:
$\hat{F}_{X}(x)=\frac{\sum_{i=1}^{n}w_{i}T_{i}I(X_{i}\leq x)}{\sum_{i=1}^{n}w_{i}T_{i}},\quad\hat{F}_{Y}(y)=\frac{\sum_{i=1}^{n}w_{i}(1-T_{i})I(Y_{i}\leq y)}{\sum_{i=1}^{n}w_{i}(1-T_{i})}$
where $I(\cdot)$ is the indicator function. It’s worth noting that linear equating can be viewed as a special case of equipercentile equating. When a kernel function (von Davier et al., , 2004; Wiberg et al., , 2024) is used to continuize the test score distributions, linear equating emerges as the limiting case of equipercentile equating as the smoothing parameter tends to infinity. Formally, if we define a smoothed version of the equipercentile equating transformation as:
$\varphi_{EP,h}(y)=F_{X,h}^{-1}(F_{Y,h}(y))$
where $F_{X,h}$ and $F_{Y,h}$ are kernel-smoothed distribution functions with bandwidth $h$, then:
@@ 4.2 Mixed-Format Tests
Our framework can be extended to handle mixed-format tests, including not only binary items but also nominal or ordinal items. Let $\mathbf{X}=(X_{1},\ldots,X_{p})$ and $\mathbf{Y}=(Y_{1},\ldots,Y_{p})$ represent the item responses for test forms $X$ and $Y$, respectively, where each item can have a different number of response categories. For nominal items, we can use multinomial logistic regression to model the propensity scores:
$\log\left(\frac{P(T_{i}=1\mid\mathbf{C}_{i})}{P(T_{i}=0\mid\mathbf{C}_{i})}\right)=\beta_{0}+\bm{\beta}^{T}\mathbf{C}_{i}$
@@
$\log\left(\frac{P(T_{i}\leq j\mid\mathbf{C}_{i})}{P(T_{i}>j\mid\mathbf{C}_{i})}\right)=\alpha_{j}-\bm{\beta}^{T}\mathbf{C}_{i},\quad j=1,\ldots,J-1$
where $J$ is the number of ordinal categories. The equating transformation for mixed format items can be defined separately for each item type and then combined to form an overall equating transformation. For example, for a test with both dichotomous and polytomous items, we might define:
@@ 5 Empirical Analysis
To empirically illustrate the proposed local equating methods and compare with existing methods, two consecutive forms from the Swedish Scholastic Assessment Test (SweSAT) were used. SweSAT is a paper and pencil test with 160 multiple-choice binary-scored items, and it consists of a quantitative section of 80 items and a verbal section of 80 items that are equated separately. The SweSAT is given twice a year and is a college admission test. Recently, an anchor test has been added but previously, different groups with specific values on their covariates have been a major part of the equating process. For details about equating methods used for the SweSAT, see Lyrén and Hambleton, (2011). The empirical study was carried out in R (R Core Team, , 2024).
The quantitative new test form X was equated to the old quantitative test form Y. To facilitate comparisons with results using a NEC design, the same data material was used as in Wiberg and Bränberg (2015). The used sample included examinees who had taken both the X and Y test forms. From these two test forms a 24-item anchor test was constructed from a selection of 12 items from each administration. Note, if an anchor test is internal or external it is irrelevant to the statistical nature of the equating procedure (van der Linden and Wiberg, , 2010; Wiberg and van der Linden, , 2011), thus this issue does not need to be addressed further here. There were 14,644 examinees who took both test forms. Two equally sized samples of 7,322 examinees for each test were used as the group was divided in half. The score distributions are displayed in Figure 1.
@@
In this analysis, we made use of the covariates gender (0 = female, 1 = male), age, and test scores from the verbal section (range: 0–80). A summary of these covariates are given in Table 1.
@@
Test Form Y correlation 0.48 --0.14 0.26 0.81
Test Form X correlation 0.52 --0.13 0.28 0.81
@@
Note. aValues in parentheses correspond to test form Y. bFor the Age variable, the statistics presented are median and quartile deviation. Age correlations are Spearman coefficients, and Gender correlations are point-biserial. All other measures are means and standard deviations.
@@ 6.1 Design
with $\theta_{i}$ representing the latent ability of examinee $i$, $a_{l}$ the discrimination parameter, and $b_{l}$ the difficulty parameter for item $l$. We simulated ability distributions according to $\theta_{P}\sim N(0,1)$ and $\theta_{Q}\sim N(0.5,1)$. Item parameters were generated to reflect realistic test characteristics: discrimination parameters $a_{l}\sim U(0.5,2)$ and difficulty parameters $b_{l}\sim N(0,1)$. The number of items and the number of anchor items were fixed at 40 and 20, respectively, and the sample size ($N$) was varied according to $N=\{1000,5000\}$. Although test forms X and Y were chosen to be smaller than the empirical illustration this conservative choice allows us to safely generalize the results to longer tests, as concluded by Wiberg and van der Linden, (2011).
The simulation included three categorical covariates with varying numbers of categories (3, 4, and 5 categories, respectively). These covariates were generated using a 2PL model structure similar to the test items, with covariate-specific discrimination parameters $a_{c}\sim U(0.5,1.5)$ and difficulty parameters $b_{c}\sim N(0,1)$. This design created a moderate correlation structure between the covariates and the latent trait. A weaker correlation structure was also investigated by generating covariate parameters with $a_{c}\sim U(0.1,0.5)$.
@@
where $\beta_{0}=0.0$, $\beta_{A}=-0.35$, $\beta_{1}=0.1$, $\beta_{2}=-0.1$, and $\beta_{3}=0.1$. The number of strata was determined through an evaluation of covariate balance, measured by the absolute standardized mean difference (ASMD) within each stratum. After examining various stratification schemes, we fixed the number of strata at 8 for all scenarios.
This design allows us to evaluate the equating methods under various realistic testing scenarios. Three equating methods were compared: the traditional anchor-based equating method, the propensity score stratification equating method, and the IPW equating method. We conducted 500 replications for each scenario considered. The simulation study was carried out in R (R Core Team, , 2024), and the code can be obtained upon request from the corresponding author.
@@ 6.2 Evaluation Measures
For each simulation replication, we sample examinees from both populations, extracting their total scores, anchor scores, covariates scores, and latent abilities. Based on the sampled data, we estimate the equating transformations for the three methods: anchor-based equating, propensity score stratification, and IPW. To evaluate bias, we partition the sample of examinees from test form $Q$ into bins based on the observed values of $\theta_{i}$, denoted as $\{\theta_{j}\}_{j=1}^{\text{nbins}}$, where each bin represents a subinterval $[\theta_{j},\theta_{j+1}]$ of the range of $\theta$. For each examinees in form $Q$ whose ability falls within the interval $[\theta_{j},\theta_{j+1}]$, the corresponding equated score $\hat{\varphi}_{M}(Y_{i})$ is computed using the equating transformation for method $M$ (anchor-based, propensity score stratification, or IPW). The equating transformations $\hat{\varphi}_{M}(Y_{i})$ transform scores from form $Q$ to the scale of form $P$ based on the specific proxy used by each method: the anchor score for anchor-based equating, the estimated propensity score strata for propensity score stratification, and the stabilized IPW weights for the IPW method.
The true equated score for an examinee in bin $[\theta_{j},\theta_{j+1}]$ is derived from the population-level distributions of $X$ and $Y$. Specifically, let $\mu_{X\mid\theta_{j}}$ and $\mu_{Y\mid\theta_{j}}$ denote the conditional means of the total scores for forms $P$ and $Q$, respectively, given that the examinee’s latent ability lies in the bin $[\theta_{j},\theta_{j+1}]$. Similarly, let $\sigma_{X\mid\theta_{j}}$ and $\sigma_{Y\mid\theta_{j}}$ denote the conditional standard deviations of the scores on forms $P$ and $Q$ given $\theta_{j}$. The true equating transformation is then:
@@
which transforms the observed score $Y_{i}$ from form $Q$ to the scale of form $P$, accounting for the differences in the conditional distributions of scores given $\theta$. To calculate the bias for each method, we compute the expected absolute difference between the equated score $\hat{\varphi}_{M}(Y_{i})$ produced by method $M$ and the true equated score $\varphi^{*}(Y_{i})$, given $\theta_{j}$. The bias for method $M$ in bin $[\theta_{j},\theta_{j+1}]$ is defined as:
@@
$\hat{\text{Bias}}_{M}(\theta_{j})=\frac{1}{\text{MC}_{\text{reps}}}\sum_{m=1}^{\text{MC}_{\text{reps}}}\frac{1}{n_{\theta_{j}}}\sum_{i\in\theta_{j}}\left|\varphi^{*}(Y_{i})-\hat{\varphi}_{M}^{(m)}(Y_{i})\right|,$
where $n_{\theta_{j}}$ is the number of examinees in bin $[\theta_{j},\theta_{j+1}]$.
In addition to bias, we evaluate the Root Mean Squared Error (RMSE) of the equating methods. The RMSE for method $M$ in bin $[\theta_{j},\theta_{j+1}]$ is defined as:
@@
$\hat{\text{RMSE}}_{M}(\theta_{j})=\sqrt{\frac{1}{\text{MC}_{\text{reps}}}\sum_{m=1}^{\text{MC}_{\text{reps}}}\frac{1}{n_{\theta_{j}}}\sum_{i\in\theta_{j}}\left(\varphi^{*}(Y_{i})-\hat{\varphi}_{M}^{(m)}(Y_{i})\right)^{2}}.$

### v2

@@ 3.2.2 Assumptions and Properties
The IPW-stratified approach offers several advantages over global weighting. First, it allows for heterogeneous equating relationships across different regions of the propensity score distribution, capturing potential effect modification by covariates. Second, by conducting weight truncation within strata, we can better control the influence of extreme weights while preserving local equating relationships. However, this flexibility comes at the cost of requiring sufficient sample sizes within each stratum to ensure stable estimation of the local equating functions.
@@ 4.1 Equipercentile Equating
While our primary focus has been on linear equating, the framework can be extended to equipercentile equating. Equipercentile equating assumes that scores on different forms of a test should have the same percentile ranks in the populations of examinees taking each test form. Let $F_{X}$ and $F_{Y}$ denote the cumulative distribution functions of scores on test forms $\mathcal{X}$ and $\mathcal{Y}$, respectively. The equipercentile equating transformation is defined as
$\varphi_{EP}(y)=F_{X}^{-1}\bigl(F_{Y}(y)\bigr).$
To implement this within our IPW framework, we estimate $F_{X}$ and $F_{Y}$ using weighted empirical distribution functions. Recall that $T_{i}=1$ for form $\mathcal{Y}$ and $T_{i}=0$ for form $\mathcal{X}$. We then define
$\hat{F}_{X}(x)=\frac{\displaystyle\sum_{i=1}^{n}w_{i}\,(1-T_{i})\,I\bigl(X_{i}\leq x\bigr)}{\displaystyle\sum_{i=1}^{n}w_{i}\,(1-T_{i})},\quad\hat{F}_{Y}(y)=\frac{\displaystyle\sum_{i=1}^{n}w_{i}\,T_{i}\,I\bigl(Y_{i}\leq y\bigr)}{\displaystyle\sum_{i=1}^{n}w_{i}\,T_{i}},$
where $I(\cdot)$ is the indicator function and $w_{i}$ are the stabilized inverse‐probability weights. It’s worth noting that linear equating can be viewed as a special case of equipercentile equating. When a kernel function (von Davier et al., 2004; Wiberg et al., 2025) is used to continuize the test score distributions, linear equating emerges as the limiting case of equipercentile equating as the smoothing parameter tends to infinity. Formally, Let $k(\cdot)$ be a symmetric, integrable kernel density with cumulative distribution function $K(u)=\int_{-\infty}^{u}k(v)\,dv$. For a bandwidth $h>0$ define the scaled (integrated) kernel $K_{h}(u)=K(u/h)$. Using the stabilized inverse-probability weights $w_{i}$ introduced earlier, we can estimate the smoothed distribution functions by
$\hat{F}_{X,h}(x)=\frac{\displaystyle\sum_{i=1}^{n}w_{i}\,(1-T_{i})\,K_{h}\!\bigl(x-X_{i}\bigr)}{\displaystyle\sum_{i=1}^{n}w_{i}\,(1-T_{i})},\qquad\hat{F}_{Y,h}(y)=\frac{\displaystyle\sum_{i=1}^{n}w_{i}\,T_{i}\,K_{h}\!\bigl(y-Y_{i}\bigr)}{\displaystyle\sum_{i=1}^{n}w_{i}\,T_{i}}.$
If we define the smoothed version of the equipercentile equating transformation as:
$\varphi_{EP,h}(y)=F_{X,h}^{-1}(F_{Y,h}(y)),$
then:
@@
This limit is taken at the population level (hence deterministic); when the population moments are replaced by their sample estimates, the same result holds in probability.
@@ 4.2 Mixed-Format Tests
Our framework can be extended to handle mixed-format tests, including not only binary items but also nominal or ordinal items. Let $\mathbf{X}=(X_{1},\ldots,X_{p})$ and $\mathbf{Y}=(Y_{1},\ldots,Y_{p})$ represent the item responses for test forms $\mathcal{X}$ and $\mathcal{Y}$, respectively, where each item can have a different number of response categories. For nominal items, we can use multinomial logistic regression to model the propensity scores:
$\log\left(\frac{P(T_{i}=1\mid\mathbf{C}_{i})}{P(T_{i}=0\mid\mathbf{C}_{i})}\right)=\beta_{0}+\boldsymbol{\beta}^{T}\mathbf{C}_{i}$
@@
$\log\left(\frac{P(T_{i}\leq j\mid\mathbf{C}_{i})}{P(T_{i}>j\mid\mathbf{C}_{i})}\right)=\alpha_{j}-\boldsymbol{\beta}^{T}\mathbf{C}_{i},\quad j=1,\ldots,J-1$
where $J$ is the number of ordinal categories, and $T_{i}$ denotes the categorical treatment assignment. The equating transformation for mixed format items can be defined separately for each item type and then combined to form an overall equating transformation. For example, for a test with both dichotomous and polytomous items, we might define:
@@ 5 Empirical Analysis
To empirically illustrate the proposed local equating methods and compare with existing methods, two consecutive forms from the Swedish Scholastic Assessment Test (SweSAT) were used. SweSAT is a paper and pencil test with 160 multiple-choice binary-scored items, and it consists of a quantitative section of 80 items and a verbal section of 80 items that are equated separately. The SweSAT is given twice a year and is a college admission test. Since 2011, an anchor test is included but previously, different groups with specific values on their covariates have been a major part of the equating process. For details about previous used equating methods for the SweSAT, see Lyrén and Hambleton (2011). Note, although anchor tests are used in the equating, covariates are still important to examine and use in the equating process to adjust for nonequivalent test taker groups. The job market highly influences which test takers that take the test and how large the test taking group is for any given year. If there is an economic recession and the unemployment rate is high, more test takers (with a diverse background) tend to take the SweSAT as opposed to when the unemployment rate is low. Furthermore, Lyrén and Hambleton (2011) found clear signs that the equivalent groups design assumptions were violated for the SweSAT. The empirical study was carried out in R (R Core Team, 2024).
The quantitative new test form $\mathcal{X}$ was equated to the old quantitative test form $\mathcal{Y}$. To facilitate comparisons with results using a NEC design, the same data material was used as in Wiberg and Bränberg (2015). The used sample included examinees who had taken both the $\mathcal{X}$ and $\mathcal{Y}$ test forms. From these two test forms a 24-item anchor test was constructed from a selection of 12 items from each administration. Note, if an anchor test is internal or external it is irrelevant to the statistical nature of the equating procedure (van der Linden and Wiberg, 2010; Wiberg and van der Linden, 2011), thus this issue does not need to be addressed further here. There were 14,644 examinees who took both test forms. Two equally sized samples of 7,322 examinees for each test were used as the group was divided in half. The score distributions are displayed in Figure 1.
@@
In this analysis, we made use of the covariates gender (0 = female, 1 = male), age, and test scores from the verbal section (range: 0–80). A summary of these covariates are given in Table 1. The choice of used covariates depended both on availability of covariates from SweSAT administrations, and the fact that these covariates have been used successfully when calculating propensity scores in previous studies, e.g. Wallin and Wiberg (2019).
@@
Test Form $\mathcal{Y}$ correlation 0.48 -0.14 0.26 0.81
Test Form $\mathcal{X}$ correlation 0.52 -0.13 0.28 0.81
@@
Note. aValues in parentheses correspond to test form $\mathcal{Y}$. bFor the Age variable, the statistics presented are median and quartile deviation. Age correlations are Spearman coefficients, and Gender correlations are point-biserial. All other measures are means and standard deviations.
@@ 5.1.3 IPW Method
5.1.4 Comparison Between the Methods
In Figure 7, we compare the anchor-based and propensity score-based equating functions within anchor-defined strata. Specifically, the sample was partitioned into three groups based on anchor scores (Low = 0–33%, Medium = 33–67%, High = 67–100%), and separate anchor-based equating functions were estimated within each stratum. We then evaluated the corresponding propensity score–based equating functions in the stratum whose median estimated propensity score fell within the same anchor tertile, so that both approaches are applied under parallel conditions. The vertical axis represents the equated score minus the anchor‐based score; the horizontal axis is the observed test score. In each panel, the line with circle markers depicts the stratified propensity‐score equating function evaluated in the stratum whose median estimated propensity score falls within the same anchor tertile; the line with triangle markers depicts the IPW equating function in that stratum. Horizontal solid lines indicate the Difference That Matters (DTM) threshold for number‐correct scoring (Dorans and Feigenbaum, 1994).
Both propensity‐score–based estimators yield nearly identical results, with deviations remaining small across the full score range in all three tertiles. In the High anchor tertile, equated scores fall within the DTM bounds for the vast majority of observed scores. By contrast, in the Low and Medium anchor tertiles, differences exceed the DTM threshold across much of the score scale, indicating that the difference between the propensity score-based and anchor score-based equating methods are the greatest when anchor difficulty is at the lower or intermediate levels. This pattern is due to the fact that examinees in the High anchor tertile are more homogeneous in both ability and covariates, so the anchor‐based and propensity score–based methods operate under nearly equivalent conditions. At lower anchor levels, greater heterogeneity remains, leading to larger discrepancies between the two approaches.
Figure 7: The difference in equated scores between the propensity score-based methods and the anchor-based method, computed within anchor-defined strata (Low, Medium, High). The horizontal solid lines represent the DTM threshold.
@@ 6.1 Design
with $\theta_{i}$ representing the latent ability of examinee $i$, $a_{l}$ the discrimination parameter, and $b_{l}$ the difficulty parameter for item $l$. We simulated ability distributions according to $\theta_{P}\sim N(0,1)$ and $\theta_{Q}\sim N(0.5,1)$. Item parameters were generated to reflect realistic test characteristics: discrimination parameters $a_{l}\sim U(0.5,2)$ and difficulty parameters $b_{l}\sim N(0,1)$. These parameter ranges are consistent with those used in prior simulation studies (see e.g., Andersson and Wiberg (2017) and reflect typical item behavior in operational tests. The total number of items was fixed at 60, with 20 common items shared between the two forms ($\mathcal{X}$ and $\mathcal{Y}$). This proportion of anchor items approximately aligns with our empirical analysis (33.33% and 31.25%, respectively). The sample size ($N$) was varied according to $N=\{1000,5000\}$. Although test forms $\mathcal{X}$ and $\mathcal{Y}$ were chosen to be smaller than the empirical illustration this conservative choice allows us to safely generalize the results to longer tests, as concluded by Wiberg and van der Linden (2011).
We included three covariates, each intended to represent observable characteristics commonly available in empirical testing contexts, such as age group, gender, or scores in other domains. These types of variables are typically reported in ordered categorical form (e.g., age bands, educational attainment levels, grouped scores), and our simulation design aims to reflect this structure. To generate these covariates in a way that captures their potential relationship with latent ability, each was constructed by summing a small number of binary indicators. For example, a covariate with four categories was generated by summing three binary variables, each simulated using a 2PL model conditional on $\theta_{i}$, with indicator-specific discrimination parameters $a_{c}\sim\text{Uniform}(0.5,1.5)$ and difficulty parameters $b_{c}\sim\mathcal{N}(0,1)$. The sum yields an integer-valued, ordinal covariate with values ranging from 0 to the number of binary components. This approach allows us to control the strength of association between each covariate and latent ability through the choice of the underlying item parameters. For comparison, we also examined a weaker correlation scenario by drawing $a_{c}\sim\text{Uniform}(0.1,0.5)$.
We note that this strategy differs from using a polytomous item response model (e.g., the graded response model). Instead of simulating a single categorical response, we model a collection of binary variables whose outcomes depend on ability, then aggregate them to obtain the covariate. This method yields observed covariates with an ordinal structure, as seen in practice, introduces a tunable correlation between the covariate and latent ability, and allows the covariates to function as observable proxies for the unmeasured confounding induced by $\theta_{i}$.
@@
where $\beta_{0}=0.0$, $\beta_{A}=-0.35$, $\beta_{1}=0.1$, $\beta_{2}=-0.1$, and $\beta_{3}=0.1$. These coefficients were deliberately chosen to achieve an approximately balanced assignment of individuals to the two test forms, with $E(p)\approx 0.5$. They also introduce moderate dependencies between the covariates, anchor scores, and form assignment, reflecting plausible selection patterns in applied testing contexts. The number of strata was determined through an evaluation of covariate balance, measured by the absolute standardized mean difference (ASMD) within each stratum. After examining various stratification schemes, we fixed the number of strata at 8 for all scenarios.
This design allows us to evaluate the equating methods under various realistic testing scenarios. We emphasize that the true latent abilities $\theta$ were used solely to define population-level equating functions and to evaluate method performance (e.g., conditional bias and RMSE), but not in the implementation of the equating methods themselves. All methods relied only on observed test scores, anchor scores, and covariates as inputs. Therefore, no scale linking was required, and no IRT-based ability estimates were used in practice.
Three equating methods were compared: the traditional anchor-based equating method, the propensity score stratification equating method, and the IPW equating method. We conducted 500 replications for each scenario considered. The simulation study was carried out in R (R Core Team, 2024), and the code can be obtained upon request from the corresponding author.
@@ 6.2 Evaluation Measures
For each simulation replication, we sample examinees from both populations, extracting their total scores, anchor scores, covariates scores, and latent abilities. Based on the sampled data, we estimate the equating transformations for the three methods: anchor-based equating, propensity score stratification, and IPW. To evaluate bias, we partition the sample of examinees from test form $\mathcal{Y}$ into bins based on the observed values of $\theta_{i}$, denoted as $\{\theta_{j}\}_{j=1}^{\text{nbins}}$, where each bin represents a subinterval $\theta_{j},\theta_{j+1}]$ of the range of $\theta$. For each examinees in test form $\mathcal{Y}$ whose ability falls within the interval $\theta_{j},\theta_{j+1}]$, the corresponding equated score $\hat{\varphi}_{M}(Y_{i})$ is computed using the equating transformation for method $M$ (anchor-based, propensity score stratification, or IPW). The equating transformations $\hat{\varphi}_{M}(Y_{i})$ transform scores from test form $\mathcal{Y}$ to the scale of test form $\mathcal{X}$ based on the specific proxy used by each method: the anchor score for anchor-based equating, the estimated propensity score strata for propensity score stratification, and the stabilized IPW weights for the IPW method.
The true equated score for an examinee in bin $\theta_{j},\theta_{j+1}]$ is derived from the population-level distributions of $X$ and $Y$. Specifically, let $\mu_{X\mid\theta_{j}}$ and $\mu_{Y\mid\theta_{j}}$ denote the conditional means of the total scores for forms $\mathcal{X}$ and $\mathcal{Y}$, respectively, given that the examinee’s latent ability lies in the bin $\theta_{j},\theta_{j+1}]$. Similarly, let $\sigma_{X\mid\theta_{j}}$ and $\sigma_{Y\mid\theta_{j}}$ denote the conditional standard deviations of the scores on test forms $\mathcal{X}$ and $\mathcal{Y}$ given $\theta_{j}$. The true equating transformation is then:
@@
which transforms the observed score $Y_{i}$ from test form $\mathcal{Y}$ to the scale of test form $\mathcal{X}$, accounting for the differences in the conditional distributions of scores given $\theta$. To calculate the bias for each method, we compute the expected absolute difference between the equated score $\hat{\varphi}_{M}(Y_{i})$ produced by method $M$ and the true equated score $\varphi^{*}(Y_{i})$, given $\theta_{j}$. The bias for method $M$ in bin $\theta_{j},\theta_{j+1}]$ is defined as:
@@
$\widehat{\text{Bias}}_{M}(\theta_{j})=\frac{1}{\text{MC}_{\text{reps}}}\sum_{m=1}^{\text{MC}_{\text{reps}}}\frac{1}{n_{\theta_{j}}}\sum_{i\in\theta_{j}}\left|\varphi^{*}(Y_{i})-\hat{\varphi}_{M}^{(m)}(Y_{i})\right|,$
where $n_{\theta_{j}}$ is the number of examinees in bin $\theta_{j},\theta_{j+1}]$.
In addition to bias, we evaluate the Root Mean Squared Error (RMSE) of the equating methods. The RMSE for method $M$ in bin $\theta_{j},\theta_{j+1}]$ is defined as:
@@
$\widehat{\text{RMSE}}_{M}(\theta_{j})=\sqrt{\frac{1}{\text{MC}_{\text{reps}}}\sum_{m=1}^{\text{MC}_{\text{reps}}}\frac{1}{n_{\theta_{j}}}\sum_{i\in\theta_{j}}\left(\varphi^{*}(Y_{i})-\hat{\varphi}_{M}^{(m)}(Y_{i})\right)^{2}}.$